OpenStudy (qqstory):

Why is the integral of cos^2(theta) d(theta) = 1/2(theta) + 1/2sin(theta)cos(theta) + C.

OpenStudy (qqstory):

\[\int\limits_{}^{} \cos ^{2} \Theta d \theta = \frac{ 1 }{ 2 }*\theta + \frac{ 1 }{ 2 } \sin \theta \cos \theta +C\]

OpenStudy (qqstory):

shouldnt the answer be 1/2(theta + sin(2theta)/2)?

OpenStudy (xapproachesinfinity):

this the intergral that you are talking about \[\int \cos^2\theta d\theta\]

OpenStudy (qqstory):

yes

OpenStudy (xapproachesinfinity):

well that's the same answer my friend

OpenStudy (xapproachesinfinity):

it is just some identities manipulation

OpenStudy (xapproachesinfinity):

both are good answers

OpenStudy (qqstory):

can you show me? how they are the same???

OpenStudy (xapproachesinfinity):

well what \[\sin 2x=?\]

OpenStudy (qqstory):

2sinxcosx

OpenStudy (xapproachesinfinity):

so ?

OpenStudy (xapproachesinfinity):

2/2 cancels right

OpenStudy (qqstory):

ya

OpenStudy (xapproachesinfinity):

then you distribute 1/2 and you get the same answer

OpenStudy (qqstory):

oh ok :D

OpenStudy (qqstory):

i am learning about trigonometric substitution right now, so does it matter which form i use?

OpenStudy (xapproachesinfinity):

no it really doesn't matter! trig integrals always come with couple of different forms lol

OpenStudy (qqstory):

ok thanks :D

OpenStudy (xapproachesinfinity):

depends on what you used to integrate! as long as the procedure is correct you shouldn't worry about the final form

OpenStudy (xapproachesinfinity):

and welcome

OpenStudy (qqstory):

Thanks for the advice

OpenStudy (xapproachesinfinity):

anytime!